# Predicting post-release defects with knowledge units (KUs) of programming languages: an empirical study

### v1

@@ 2.3 Detection of KUs
Analogously to our prior work (Ahasanuzzaman et al., 2024), we employ static analysis to detect KUs from a given Java project release. First, we parse all Java source files of the release using the Eclipse JDT framework (Eclipse, 2020). The JDT framework creates an Abstract Syntax Tree (AST) and provides visitors for every element of such tree, including names, types (class, interfaces), expressions, statements, and declarations. [^3] The grammar employed by JDT can be seen at https://github.com/eclipse-jdt/eclipse.jdt.core/blob/master/org.eclipse.jdt.core.compiler.batch/grammar/java.g
@@ 3 Data Collection
Our data collection process contains five main steps. First, we collect the defect dataset. Then, we retrieve the source code for the studied projects and releases. Next, we quantify the incidence of KUs in the source code of our studied projects. We also extract traditional code metrics from the source code of our studied releases. Finally, we create KU-based feature dataset and code metric dataset. Figure 2 presents an overview of our data collection process. We describe each step in more detail below.
@@
  • (Step 1) Collecting defect dataset.
    • (Step 1.1) Download the defect dataset. We choose the defect dataset that was curated by Yatish et al. (2019). Instead of relying on heuristic to identify post-release defects, the authors employ an approach that “leverages the earliest affected release that is realistically estimated by a software development team for a given defect”. We download the defect dataset from the article’s replication package [^4] https://github.com/awsm-research/replication-icse2019. The dataset contains post-release defect data for 28 releases that span a total of 8 Java project. Table 2 presents a statistical summary of the dataset.
@@
  • (Step 2) Collecting the source code data of the studied releases.
    • (Step 2.1) Download the source code of the studied releases. To download the source code of the studied software projects, we clone the repository of each project using the command git clone project_url where project_url is the url of the repository of the project in GitHub. Then, we extract the source code of a particular release using the command git checkout tags/tag_name. Here, tag_name is the name of one of the studied release for the project. The list of projects and their associated GitHub repository links can be found in our supplementary material.
  • (Step 3) Collecting KUs from the studied releases.
    • (Step 3.1) Determining the incidence of KUs in the source code. We count the occurrences of KUs in each Java file for every studied release in the defect dataset using our custom detector, which is built on top of Eclipse JDT (refer to Section 2.3).
  • (Step 4) Collecting code metrics from the studied releases.
@@
    • (Step 4.1) Extract traditional code metrics from source code. In our study, we compare the incidence of KUs with traditional code metrics. To extract the code metrics we use the Understand tool [^5] https://scitools.com. We run the tool on the source code of each studied release and extract metrics for each Java file in the defect dataset. We extract 54 product metrics. Examples of product metrics: CountLineCode (number of lines of code), CountDeclMethodPublic (number of declared public methods in a class), and CountClassCoupled (number of other classes to which a class is coupled to). We also collect 5 process metrics: the number of commits (COMM), the number of lines added (ADDED_Lines), the number of lines deleted (DEL_LINES), the number of active developers (ADDEV) and the number of distinct developers (DDEV). Therefore, we extract a total of 59 metrics. Prior studies also use the subset of our extracted code metrics in their studies (Yatish et al., 2019; Nuñez-Varela et al., 2017; Jiarpakdee et al., 2020; Esteves et al., 2020). The full list of the collected 59 metrics for our study can be seen in B.
  • (Step 5) Creating the KU-based feature dataset and the code metrics dataset.
    • (Step 4.1) Vector representation of every file of the studied releases. Each file $F$ of a studied release $R$ is represented by two vectors in the dataset: one for KU-based features and one for code metrics. The vector for KU-based features has a length of 28, representing all the 28 identified KUs (See Table 8 in the Appendix). Each position in the vector stores the count value of a KU that is extracted from $F$. The code metrics vector has a length of 59, representing all the traditional code metrics that we study. Each position in the vector stores the value of a code metric that is extracted from $F$.
      The set of all KU vectors of all files of the studied releases can be seen as a table (dataset). We refer to it as the KU-based feature dataset ($D_{ku}$). Analogously, we refer to the set of all code metric vectors of all files of the studied releases as the code metrics dataset ($D_{cm}$). $Dataset(R,D_{ku})$ returns the KU-based feature dataset of a release $R$ and $Dataset(P,D_{cm})$ returns the code metrics dataset of the release $R$.
Summary Defect dataset: The defect dataset that was curated by Yatish et al. (2019). Data source for the source code: GitHub Pieces of data collected: Snapshots of the source code of the studied projects releases Datasets produced: KU-based feature dataset and traditional code metrics dataset
4 Preliminary Study: Do KUs offer a complementary perspective to the one offered by traditional code metrics?
Motivation. We compare KUs with traditional code metrics which are also derived from the same raw data (i.e., source code) of software systems. We conjecture that KUs provide a new perspective for the analysis of software code compared to traditional code metrics.
Approach. Our preliminary study is conducted in three steps, which are briefly explained below.
  • Step 1) Comparing KUs to code metrics via correlation analysis. As a first step to determine how KUs and code metrics differ from one another, we calculate the Spearman’s rank correlation coefficient ($rho$) between them. More formally, we calculate this correlation between all pairs of variables $x$ and $y$, where $x$ is a column in $D_{ku}$ and y is a column in $D_{cm}$. We choose the Spearman’s $\rho$ in lieu of other correlation coefficients (e.g., Pearson’s correlation) because the Spearman’s $\rho$ does not require normally distributed data. We assess Spearman’s $\rho$ using the following thresholds (Evans, 1996), which operate on 2 decimal places only: very weak for $|\rho|\leq 0.19$, weak for $0.20\leq|\rho|\leq 0.39$, moderate for $0.40\leq|\rho|\leq 0.59$, strong for $0.60\leq|\rho|\leq 0.79$, and large for $|\rho|\geq 0.80$.
  • Step 2) Clustering the files of the studied releases according to KUs and code metrics. We first cluster the files of a project using KUs. Next, we re-cluster the files of the project using code metrics and evaluate how the two clusterings differ. We use the k-means clustering algorithm (Comaniciu and Meer, 2002) to perform the clustering. K-means is one of the most popular clustering algorithms and has been widely used in software engineering research (Bishnu and Bhattacherjee, 2012; Yoon et al., 2007; Jureczko and Madeyski, 2010; Zhang et al., 2016). The detailed sub-steps are shown below:
    • Step 2a) Reducing the dimensions of the datasets. Many clustering algorithms (e.g., k-means algorithm) have inferior performance for high dimensional data  (Parsons et al., 2004). To overcome the dimensionality issue, we reduce the dimension of our data using PCA before running the k-means algorithm. We perform the PCA on both KUs dataset $(Dataset(P,D_{ku}))$ and code metric dataset $(Dataset(P,D_{cm}))$ of a project $P$ (see Section 3) using the PCA class of scikit-learn package [^6] https://scikit-learn.org/dev/modules/generated/sklearn.decomposition.PCA.html. We reduce the dimension of our data by selecting the first-n PCA components that account for at least 90% of the variance of the data.
    • Step 2b) Identify optimal number of clusters. The k-means algorithm takes the number of clusters as an input parameter (K). To identify the optimal number of clusters, we analyze the silhouette coefficient (Rousseeuw, 1987). The silhouette coefficient is a popular measure that has been widely used in prior studies (Nidheesh et al., 2020; Al-Subaihin et al., 2016; Panichella et al., 2013) for determining the optimal number of clusters. The silhouette coefficient ranges from -1.0 to 1.0. A coefficient close to one indicates that the object well-matches its own cluster and, simultaneously, poorly-matches its neighboring clusters. If most objects have a high coefficient, then the clustering configuration is deemed appropriate. We run k-means for K values ranging from 2 to 40 and calculate the mean silhouette for each K value. We determine the optimal number of clusters by selecting the K that maximizes the mean silhouette score.
    • Step 2c) Clustering the files of the studied releases. To cluster the files of the studied releases using the k-means algorithm, we employ the KMeans class of scikit-learn from python [^7] https://scikit-learn.org/1.5/modules/generated/sklearn.cluster.KMeans.html. We use the selected PCA components as the input dataset. Finally, we cluster the studied releases using the k-means algorithm and visualize the clusters using the fviz_cluster function of the factoextra R package (Kassambara and Mundt, 2017).
  • Step 3) Comparing the two clusterings. We compare the two clusterings produced in the prior step. This comparison is done both visually and then statistically through the computation of the adjusted Rand index. The Rand index refers to the probability that the two clusterings will agree on a randomly chosen pair of data points. The adjusted Rand index employs a correction-for-chance to the original Rand index. Intuitively, the adjusted Rand index is a measure of the similarity between two clusterings. The index ranges from zero (i.e., the two clusterings do not agree on any pair of points) to one (i.e., the clusterings are exactly the same).
Findings. Observation 1) The majority of KUs do not exhibit a strong correlation with code metrics. Table 3 presents the strength of correlation between KUs and code metrics. We observe that no KUs show a large correlation with code metrics. We also observe that only 5.6% of the identified KUs (e.g., Operator and Decision KU) show a strong a correlation with code metrics. However, the Spearman’s $\rho$ for all pairs consisting of a code metric and a KUs is below 0.60 for the majority of KUs (i.e., 94.4% of the identified KUs). Therefore, the majority of KUs do not exhibit a strong correlation with code metrics. This result indicates that KUs can provide an additional perspective for the analysis of software code.
Table 3: The strength of correlation between KUs and code metrics.
Strength of correlation The percentage of KUs
Large correlation ($|\rho|\geq 0.80$) 0%
Strong correlation ($0.60\leq|\rho|\leq 0.79$) 5.6%
Moderate correlation ($0.40\leq|\rho|\leq 0.59$) 22.2%
Weak correlation ($0.20\leq|\rho|\leq 0.39$) 33.3%
Very weak correlation ($|\rho|\leq 0.19$) 38.9%
[…]
Figure 3: Identified clusters for knowledge units (KUs) and code metrics of the “Wicket-1.3.0-beta2” release.
Observation 2) KUs identify two times more clusters than code metrics. We observe that the clustering based on KUs produce 4 (median) clusters, whereas the clustering of code metrics produce e (median) clusters. In other words, the Java files of the studied releases can be seen as belonging to two times more clusters through the lens of KUs than the clusters that are identified by code metrics. Figure 3 depicts the identified clusters for KUs and code metrics for the “Wicket-1.3.0-beta2” [^8] https://github.com/structurizr/java. We observe that the clustering based on the lens of KUs produce 4 clusters and code metrics produce 2 clusters.
Finally, we observe that the median value of adjusted Rand index between the two clusterings across the studied releases is 0.54 which is lower than 1. In other words, the probability that the two clusterings will agree on a randomly chosen studied release is 54% (after correction-for-chance adjustment). Since a Rand index value other than 1 indicates differences in clusters, we conclude that the clusters of KUs are not the same as those of code metrics.
Figure 4: Visualization of clusters for KUs and code metrics. Each polygon represents a cluster of KUs and the dotted points of colors represent files in the clusters of code metrics.
Observation 3) The identified clusters via code metrics do not completely overlap with the identified clusters via KUs. Figure 4 shows the overlap between the two clusterings of the “Wicket-1.3.0-beta2” project. Each polygon represents a cluster of KU and each colored point represents a file in one of the clusters of the code metric clustering. As shown in Figure 4, cluster 2, 3 and 4 of KUs (i.e., the polygons) contain files of different clusters of code metrics (i.e., each polygon contains different coloured points). For example, cluster 4 of KUs contains 33.3% files of cluster 0 of code metrics and 67.7% files of cluster 1 of code metrics. Therefore, these KU clusters are not completely overlapped by a single cluster of code metrics (i.e., a KU cluster contains files of more than one cluster of code metrics). This result is also an indication that KUs offer a complementary perspective for analyzing software code.
Figure 5: The distribution of the percentage of KU clusters that are not overlapped by a single cluster of code metrics.
Figure 5 presents the distribution of the percentage of KU clusters that are not overlapped by a single cluster of code metrics across the studied releases. The median of the percentage of KU clusters that are not overlapped by a single code metric cluster is 81.67%, indicating that code clusters do not completely overlap with the KU clusters (i.e., polygons have a somewhat mix of different colored dots). Such a result indicates that there exists KU clusters that are not uniquely described by a single cluster of code metrics. This result is also an indication that KU clusters can capture a different perspective of software systems from the identified clusters via code metrics.
Summary Preliminary Study: Do KUs offer a complementary perspective to the one offered by traditional code metrics? KUs offer a new lens through which the code of software systems can be analyzed, revealing a perspective that is complementary to that yielded by traditional code metrics. In particular: None of the KUs exhibit a large correlation with code metrics. Through the lens of KUs, the files of the studied projects could be grouped into two times more cluster than the identified clusters via code metrics. The identified clusters using code metrics do not completely overlap with the identified clusters using KUs.
@@ 5 Findings
In this section, we address our research questions. For each question, we discuss our motivation for studying it, the approach that used to answer it, and the findings that we observe.
5.1 RQ1: How effective are KU-based features for identifying defect prone code?
5.1.1 Motivation
@@
From our preliminary study, we conclude that KUs offer a new lens through which the code of software systems can be analyzed. This conclusion motivates us to study the effectiveness of KUs in post-release defect prediction problem. Traditionally, defect prediction has relied heavily on code metrics, such as process and product metrics (Gyimothy et al., 2005; Moser et al., 2008; Zhang, 2009; Shihab et al., 2010). We believe that exploring novel types of data for feature engineering is as important as developing or assessing new machine learning approaches for defect prediction. The more indicators researchers have, the more ammunition they have to understand bugs and conjecture their real causes.
5.1.2 Approach
@@
First, we build a defect prediction model using the KU-based features for each of the studied releases. We refer to this model as KUCLS. Next, we evaluate the performance of the KUCLS. Finally, we compare the performance of the KUCLS to those models that are built using traditional code metrics. In the following, we explain each of the steps that we follow during this process.
Figure 6: An overview of building the KUCLS.
@@
  • (Step 1) Construct the KUCLS model. In this step, we construct and evaluate KUCLS prediction model (for clean/buggy classes). An overview of our approach for building the KUCLS can be seen in Figure 6.
@@
    • (Step 1.1) Select non-correlated features. Prior to building any defect prediction model, it is recommended to mitigate correlated features (Jiarpakdee et al., 2019) to better interpret the results of a model. Therefore, we want to check the performance of our studied models using non-correlated features. We select non-correlated features using the approach proposed by Jiarpakdee et al. (2018). This approach is provided by the AutoSpearman function in the python implementation [^9] https://xai4se.github.io/defect-prediction/data-preprocessing.html. AutoSpearman automatically selects non-correlated metrics based on two analyses: (1) a Spearman rank correlation test and (2) a Variance Inflation Factor analysis. We apply the AutoSpearman on KU-based feature dataset, which resulted in 28 non-correlated KU-based features (i.e., no KU-based feature is removed).
@@
    • (Step 1.2) Generate bootstrap samples. To estimate the performance of our models in practice, we employ the out-of-sample bootstrap validation technique with 100 repetitions (Tantithamthavorn et al., 2017; Efron, 1983). In this technique, we randomly sample dataset with replacement to create 100 bootstrap samples. Each bootstrap sample have the same number of data points as the original dataset, but some points may appear multiple times while others may be excluded. These excluded data points from each bootstrap sample form the corresponding out-of-sample data, which we use for testing our models. We train our models on each bootstrap sample and test its performance on the associated out-of-sample data.
    • (Step 1.3) Construct the KUCLS model. We use the non-correlated KU-based features to construct our KUCLS model. In terms of the machine learning algorithm, we choose a random forest classifier with default parameters values (scikit-learn python implementation). This parameter setting of random forest classifier is identified as the optimal choice for achieving the best performance in classifying post-release defects (Yatish et al., 2019).
@@
    • (Step 1.4) Evaluate the KUCLS model. To evaluate the accuracy of our defect prediction models, we calculate the Area Under the receiver operator characteristic Curve (AUC) as it is a robust, threshold-independent measure (Tantithamthavorn and Hassan, 2018). AUC value ranges from 0 to 1, with 0.5 being as good as random guessing. The AUC values closer to one indicate a model that is good at distinguishing the class of interest (buggy) from the other class (clean). Values closer to zero indicate a model that is good at predicting the class that is not of interest (clean). Since we are employing an out-of-sample bootstrap model validation with 100 repetitions, we obtain 100 AUC values for each prediction model that we evaluate.
  • (Step 2) Construct baseline models. Our main goal in RQ1 is to understand the efficacy of KUs in predicting software defects. Ultimately, we wish to understand whether we can obtain richer results (e.g., more accurate predictions) when using KUs in lieu of baselines. To keep things fair and more straightforward, we compare features that are derived from the same raw data (i.e., source code). Since KUs are derived from the source code, we compare KUs with traditional code metrics which are also derived from the source code. Moreover, previous studies widely used these subsets of our extracted code metrics for building and analyzing defect prediction models (Zhang et al., 2016; Yatish et al., 2019; Nuñez-Varela et al., 2017; Jiarpakdee et al., 2020; Laaber et al., 2021; Esteves et al., 2020). Therefore, we use the traditional code metrics and construct two baseline models: (i) a model built with product code metrics (CC_PROD), and (ii) a model built with both product and process metrics (CC). These baseline models are constructed following the procedure described in Step 1.
@@
  • (Step-3) Compare models’ performance. To check whether the performance difference between a given pair of models (e.g., KUCLS vs CC_PROD) is statistically significant, we apply the Wilcoxon signed-rank test. The Wilcoxon signed-rank test is a non-parametric statistical test for paired data. We infer that there is a statistically significant difference between the two input distributions when the p-value computed by the test is smaller than 0.05 (i.e., $alpha=0.05$). In addition, we also use the Cliff’s delta ($d$) effect size measure to quantify the practical difference between the two distributions (Long et al., 2003). We use the following thresholds for interpreting $d$ (Romano et al., 2006): negligible for $|\delta|\leq 0.147$, small for $0.147<|\delta|\leq 0.33$, medium for $0.33<|\delta|\leq 0.474$, and large otherwise.
@@
    To evaluate a model’s performance improvement compared to a baseline, we calculate the normalized AUC improvement metric. This metric measures the extent of improvement achieved by the model relative to the maximum possible improvement. It is calculated using the following formula:

### v2

@@ 2.3 Detection of KUs
Analogously to our prior work (Ahasanuzzaman et al., 2024), we employ static analysis to detect KUs from a given Java project’s release. First, we parse all Java source files of the release using the Eclipse JDT framework (Eclipse, 2020). The JDT framework creates an Abstract Syntax Tree (AST) and provides visitors for every element of such tree, including names, types (class, interfaces), expressions, statements, and declarations. [^2] The grammar employed by JDT can be seen at https://github.com/eclipse-jdt/eclipse.jdt.core/blob/master/org.eclipse.jdt.core.compiler.batch/grammar/java.g
@@ 3 Data Collection
Our data collection process is straightforward and contains four steps (Figure 2). We collect the defect dataset from the replication package of Yatish et al. (2019) (Step 1) and filter-in traditional metrics (Step 2). Then, we download the source code of our studied releases (Step 3) and compute the per-file KU metrics using our custom KU detector (Step 4). In the following, we describe each step of our data collection process in more detail.
@@
  • (Step 1) Download the defect dataset. We choose the defect dataset that was curated by Yatish et al. (2019). Instead of relying on heuristics to identify post-release defects, the authors employ an approach that “leverages the earliest affected release that is realistically estimated by a software development team for a given defect”. We download the defect dataset from the article’s replication package [^3] https://github.com/awsm-research/replication-icse2019. The dataset contains post-release defect data for 28 releases that span a total of 8 Java project. Table 2 presents a statistical summary of the dataset (adapted from original paper).
@@
  • (Step 2) Filter-in traditional metrics (product, process, and ownership) We filter-in the traditional metrics from the defect dataset. This filtered dataset contains 54 product, 5 process, and 6 code ownership metrics (65 total) calculated for each Java source code file. Product metrics were calculated using the Understand tool (SciTools, 2000). Examples include the number of lines of code (CountLineCode), number of declared public methods in a class (CountDeclMethodPublic), and number of other classes to which a class is coupled to (CountClassCoupled). The process metrics are: the number of commits (COMM), the number of lines added (ADDED_LINES), the number of lines deleted (DEL_LINES), the number of active developers (ADDEV) and the number of distinct developers (DDEV). The ownership metrics are: number of minor authors (developers contributing less than 5% of the code in a file), number of major authors (developers contributing 5% or more of the code in a module), and ownership ratio (the proportion of lines of code written by the developer who has the highest contribution of lines of code on the module). These ownership metrics are calculated for two granularities of code changes: lines of code (MINOR_LINE, MAJOR_LINE, and OWN_LINE) and commit (MINOR_COMMIT, MAJOR_COMMIT, and OWN_COMMIT).
    We refer to these 65 metrics as traditional because they have frequently been used in the literature to build and analyze defect prediction models (Zhang et al., 2016; Nuñez-Varela et al., 2017; Jiarpakdee et al., 2020; Laaber et al., 2021; Esteves et al., 2020). The full list of the traditional metrics can be seen in B.
  • (Step 3) Download the source code of the studied releases. To download the source code of the studied software projects, we clone the repository of each project using the command git clone project_url where project_url is the url of the repository of the project in GitHub. Then, we extract the source code of a particular release using the command git checkout tags/tag_name. Here, tag_name is the name of one of the studied release for the project.
  • (Step 4) Compute KU metrics. We count the occurrences of each KU in each Java file from every studied release in the defect dataset using our custom KU detector (refer to Section 2.3 for details).
The outputs of the data collection process are the KU metrics dataset and the traditional metrics dataset (Figure 2). These datasets are employed to build several defect prediction models as part of our RQs.
@@ 4 Findings
In this section, we address our research questions. For each question, we discuss our motivation for studying it, the approach that we used to answer it, and our findings.
4.1 RQ1: How effective are KU features for predicting post-release defects?
@@
Motivation. Traditionally, defect prediction has relied heavily on process, product, and code ownership metrics (Gyimothy et al., 2005; Moser et al., 2008; Zhang, 2009; Shihab et al., 2010) (traditional metrics). KUs offer a new lens through which the code of software systems can be analyzed, since they capture system traits that are not captured by those traditional metrics. More generally, we believe that exploring novel types of data for feature engineering is as important as developing or assessing new machine learning approaches for defect prediction. The more indicators researchers have, the better equipped they are to understand defects and hypotheisze their underlying causes.
@@
Approach. First, we build a defect prediction model using the KU metrics dataset for each of the studied releases. We treat each KU metric as a model feature and we refer to the entire set of KU features as the KU feature set. Also, we refer to this KU model as KUM . Next, we evaluate the performance of KUM based on an out-of-sample bootstrap model validation technique. Finally, we compare the performance of KUM to baseline models that are built using the traditional metrics dataset. In the following, we explain our approach in more detail:
Figure 3: An overview of our approach for building and evaluating KUM.
@@
  • (Step 1) Construct KUM. In this step, we construct and evaluate KUM. KUM is a binary classification model that outputs either clean or defective for an input Java file. An overview of our approach for building KUM can be seen in Figure 3.
@@
    • (Step 1.1) Select non-correlated features. Prior to building any defect prediction model, it is recommended to mitigate correlated features (Jiarpakdee et al., 2019) to better interpret the results of a model. Therefore, we want to check the performance of our studied models using non-correlated features. We select non-correlated features using AutoSpearman (Jiarpakdee et al., 2018). AutoSpearman automatically selects non-correlated features based on two analyses: (1) a Spearman rank correlation test and (2) a Variance Inflation Factor (VIF) analysis. Applying AutoSpearman [^4] https://xai4se.github.io/defect-prediction/data-preprocessing.html to the KU feature set resulted in the selection of all features, as no feature correlations were detected.
@@
    • (Step 1.2) Generate bootstrap samples. To estimate the performance of our models in practice, we employ the out-of-sample bootstrap validation technique with 100 repetitions (Tantithamthavorn et al., 2017; Efron, 1983). In this technique, we randomly sample dataset with replacement to create 100 bootstrap samples. Each bootstrap sample has the same number of data points as the original dataset, but some points may appear multiple times while others may be excluded. These excluded data points from each bootstrap sample form the corresponding out-of-sample data, which we use for testing our models. We train our models on each bootstrap sample and test its performance on the associated out-of-sample data.
    • (Step 1.3) Construct KUM. We use the KU feature set to construct KUM . Similarly to Yatish et al. (2019), we use a random forest classifier with default parameters values (scikit-learn python implementation). For a detailed analysis of the influence of classifier choice and hyperparameter tuning on model performance, please refer to C.
@@
    • (Step 1.4) Evaluate KUM. To evaluate the accuracy of our defect prediction models, we calculate the Area Under the receiver operator characteristic Curve (AUC). AUC is a robust, threshold-independent measure (Tantithamthavorn and Hassan, 2018). AUC value ranges from 0 to 1, with 0.5 being as good as random guessing. The AUC values closer to one indicate a model that is good at distinguishing the class of interest (defective) from the other class (clean). Values closer to zero indicate a model that is performing worse than random guessing, consistently misclassifying instances by predicting the opposite class (i.e., labeling defective instances as clean and vice versa). Since we are employing an out-of-sample bootstrap model validation with 100 repetitions, we obtain 100 AUC values for each prediction model that we evaluate.
  • (Step 2) Construct baseline models. Our main goal in RQ1 is to understand the efficacy of KUs in predicting software defects. To construct baseline models, we use the traditional metrics dataset that we extracted from our data collection. We construct four baseline model: (i) a model built with product metrics (PROD), (ii) a model built with process metrics (PROC), (iii) a model built with code ownership metrics (OWN) and (iv) a model built with product, process, and ownership metrics (i.e., all the traditional metrics – TM). These baseline models are constructed following the same procedure described in Step 1.
@@
  • (Step-3) Compare models’ performance. To check whether the performance difference between a given pair of models (e.g., KUM vs PROD) is statistically significant, we apply the Wilcoxon signed-rank test. The Wilcoxon signed-rank test is a non-parametric statistical test for paired data. We infer that there is a statistically significant difference between the two input distributions when the p-value computed by the test is smaller than 0.05 (i.e., $alpha=0.05$). In addition, we also use the Cliff’s delta ($d$) effect size measure to quantify the practical difference between the two distributions (Long et al., 2003). We use the following thresholds for interpreting $d$ (Romano et al., 2006): negligible for $|\delta|\leq 0.147$, small for $0.147<|\delta|\leq 0.33$, medium for $0.33<|\delta|\leq 0.474$, and large otherwise.
@@
    To evaluate a model’s performance improvement compared to a baseline, we define a metric called normalized AUC improvement. This metric measures the extent of improvement achieved by the model relative to the maximum possible improvement. It is calculated using the following formula: